OpenStudy (anonymous):

Given segment XZ and point Y that lies on XZ, find ZY if XZ=17, XY = -11+5x, and ZY = 33-6x

OpenStudy (anonymous):

XY and ZY added together form XZ (The whole segment) XY+ZY=XZ Substitute the values in to get: -11+5x+33-6x=17 Combine like terms to get: -x+22=17 Isolate x by subtracting 22 from 17 You end up with -x=-5 Divide both sides by 1 to get: x=5 Substitute the x in 33-6x with 5 You will have 33-6(5)=ZY 33-30=ZY ZY=3

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Glad to be of service.